SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Trigger Operation Meeting (03/04/08)

Attendees: Su Dong, Rahmat Rahmat

Meeting Information

 Date      :  March 4th, 2008
 Time      :  2:00pm - 3:30pm PST
 SLAC Room :  Madrone conf room, ROB 
 
 Dial-in   :  1-510-665-5437, pass code 6649#

--- Minutes ---

Operations


    We started running Y(2S) on March 1, 2008

    The peak of TSF SL1 was shifted, probably due to background,

    On Sunday there was a new dip in TSF SL 6 run 80119, but after the reconfiguration the dip disappeared
    (Su Dong & Rahmat called IR-2)

    There was no beam due to vacuum leak since Monday and still not taking (colliding) data till Tues evening

    Hojeong
    Su Dong asked me to check if there was any IFR/IFT related work during Run range 79253->79285. I checked with subsystem managers especially with IFR commissioner if there was ANY activity including power cycle, ( I let him know we are particularly interested in forward end electronics but I explicitly asked him to look for ALL activities) and he didn't find anything unusual.

EMT

    No failed Calibration

DQM

    Rahmat is checking
      Y(3S) runs
      Y(2S) onpeak runs
      Y(2S) scan runs

Trigger Configuration


This page was created by Rahmat; however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.

Last Modified: Feb 26th, 2008.